OpenStudy (anonymous):

square root of 52 simplifying expressions with complex numbers

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{52} = \sqrt{4*13}\] \[\large \sqrt{52} = \sqrt{4}*\sqrt{13}\] \[\large \sqrt{52} = ???\]

OpenStudy (anonymous):

is that the answer

jimthompson5910 (jim_thompson5910):

no the final answer is something you need to find

jimthompson5910 (jim_thompson5910):

ie you need to replace the ??? with the answer

jimthompson5910 (jim_thompson5910):

what is the square root of 4?

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

so... \[\large \sqrt{52} = \sqrt{4*13}\] \[\large \sqrt{52} = \sqrt{4}*\sqrt{13}\] \[\large \sqrt{52} = 2\sqrt{13}\]

jimthompson5910 (jim_thompson5910):

now you have the final answer

OpenStudy (anonymous):

thanks you are going to be my new math tutor

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

are you going to

jimthompson5910 (jim_thompson5910):

sure I can help out, what exactly do you have in mind when you say that though?

OpenStudy (anonymous):

about you being my math tutor?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

if i need help with math

jimthompson5910 (jim_thompson5910):

alright I can help out from time to time also there are lots of other helpers on this site (and not just me)

OpenStudy (anonymous):

so you dont want to help me

jimthompson5910 (jim_thompson5910):

I do and I can I'm just saying I'm not the only one who can